LASER SPECTROSCOPY OF CoI AND NiI

\underline{J. W-H. LEUNG}, A. L. WONG, W. S. TAM, A. S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong; SHUIMING HU, Department of Chemical Physics, University of Science and Technology of China, Hefei, Anhui, China.